You need to prepare an acetate buffer of pH 5.77 f
Answer #1

no of moles of CH3COOH = MOlarity* volume in L

                                         = 0.671*0.625   = 0.4194 moles

PH   = PKa + log[CH3COOK]/[CH3COOH]

5.77 = 4.76 + log[CH3COOK]/0.4194

5.77-4.76    = log[CH3COOK]/0.4194

1.01      = log[CH3COOK]/0.4194

[CH3COOK]/0.4194 = 101.01

[CH3COOK]/0.4194 = 10.23

[CH3COOK]              = 10.23*0.4194

[CH3COOK]              = 4.29

no of moles of CH3COOK = 4.29moles

no of moles of CH3COOK   = no of moles of KOH

no of moles of KOH         = 4.29 moles

volume of KOH               = no of moles /molarity

                                       = 4.29/2.12 = 2.023L = 2023ml = 2.023*103ml>>>> answer